OpenStudy (anonymous):

-3sin(0.5piN)+4cos(0.5piN)= 5cos(0.5piN+0.6435) ..how i can convert from left to right side?

OpenStudy (anonymous):

So we want \[-\sin(C)\sin(0.5\pi N)+\cos(C)\cos(0.5 \pi N)\] So we can write this as \[\cos(0.5 \pi N+C)\] \[\frac{\sqrt{(3)^2+(4)^2}}{\sqrt{(3)^2+(4)^2}}(-3 \sin(0.5 \pi N)+4 \cos(0.5 \pi N))\] \[5(\frac{-3}{5}\sin(0.5 \pi N)+\frac{4}{5} \cos(0.5 \pi N))\]

OpenStudy (anonymous):

But we want sin(C)=3/5 and cos(C)=4/5

OpenStudy (anonymous):

which can happen in a (3,4,5) triangle :)

OpenStudy (anonymous):

\[5\cos(0.5 \pi N+C)=-3 \sin(0.5 \pi N)+4 \cos(0.5 \pi N)\]

OpenStudy (anonymous):

Divide by 5 on the LHS first

OpenStudy (anonymous):



OpenStudy (anonymous):

You need to find angle C between 0 and pi/2

OpenStudy (anonymous):

you can use either of those equations I mentioned

OpenStudy (anonymous):

Then use @ZombiePig 's diagram express 3/5 and 4.5 in terms of cos C and sin C

OpenStudy (anonymous):

*4/5

OpenStudy (anonymous):

Then apply cos a . cos b - sin a . sin b = cos (a+b)

OpenStudy (anonymous):

Can this be solved without electronic aid? (Don't say any chart or table)

OpenStudy (anonymous):

yep unless you want to approximate C

OpenStudy (anonymous):

But still we can guesstimate a value between 0 and pi/2 without a calculator

OpenStudy (anonymous):

I got it ! thanks all :)

OpenStudy (anonymous):

\[C=\arccos(\frac{4}{5}) \text{ exact value}\]

OpenStudy (anonymous):

\[{-3 \over 5}\sin(0.5piN)+{4 \over 5}\cos(0.5piN)\] \[{-\sin C}\sin(0.5piN)+{\cos C}\cos(0.5piN)\] \[\cos ( 0.5piN + C)\] \[\cos ( 0.5piN + \tan^{-1} { 3 \over 4})\]

OpenStudy (anonymous):

The question ask for proof.

OpenStudy (anonymous):

@FoolForMath , didn't get what you are trying to convey ?

OpenStudy (anonymous):

0.6435 that's C isn't? can we compute this value without electronic aid?

OpenStudy (anonymous):

I don't think so. @FoolForMath

OpenStudy (anonymous):

the question is not asking for the proof , it's just simplifying the answer.